OpenStudy (anonymous):

Please help! What is the value of the fourth term in a geometric sequence for which a1 = 30 and r = 1/2?

OpenStudy (ranga):

Without using the formula for the nth term you can find the 4th term easily: Geometric sequence is formed by multiplying each term by the common ratio r. 1st term = 30 2nd term = 30 * 1/2 = 15 3rd term = 15 * 1/2 = 7.5 4th term = ?

OpenStudy (anonymous):

how would i find the 4th term? im confused with how the 1st and 2nd term went from 30 to the 3rd term being 15

OpenStudy (ranga):

The current term = previous term * r

OpenStudy (ranga):

2nd term = 1st term * 1/2 3rd term = 2nd term * 1/2 4th term = 3rd term * 1/2

OpenStudy (anonymous):

so then 4th term would be 7.5*1/2=3.75 ?

OpenStudy (ranga):

yes.

OpenStudy (ranga):

You can also leave it as a fraction. 30, 15, 15/2, 15/4, 15/8, 15/16, ... will be the geometric sequence. The fourth term is all they want and so it is 15/4

OpenStudy (anonymous):

oh okay thanks! :)

OpenStudy (ranga):

welcome.